# Heterogeneous Graph Neural Networks for Data-driven Traffic Assignment

### v1

@@ 4.2 Experiments on Generalized Synthetic Networks
In this set of experiments, we study the performance of the proposed model when generalized to networks with different topologies. To do so, we generate synthetic networks by starting with a grid graph and adding links between randomly selected nodes in the grid. Then, a few nodes and edges are randomly removed to emulate the real road network and increase the variability. It should be noted that the proposed HetGAT involves a node feature whose dimension is equal to the total number of nodes in the graph. Therefore, our model can be generalized to different graphs as long as the total number of nodes remain the same. Therefore, the various topologies in synthetic road networks all have the same number of nodes. Two sizes of the synthetic generalized dataset are included in the experiments: 100 and 300. For each graph size, 100 different graph topologies are generated. Three of these randomly generated graphs are shown in Figure 5. The OD demand and the link capacity are also randomly generated using the scaling factor according to Equation 16 and 17, respectively. Furthermore, the UE-based traffic assignment is used to generate the ground truth data, similarly to Section 4.1.
Figure 5: The illustrations of sampled generalized synthetic networks. The network size is 100. The link color represents the link capacity of each link.
It should be noted that in fully connected neural networks (FCNN), the output dimension and the order of the entries should be fixed. This is while in our synthetic networks, even though the number of nodes are fixed, the number of edges are not necessarily the same in different topologies. Because of this, FCNN is not suitable for experiments on generalized synthetic networks. Therefore, GAT and GCN are the only two models that are compared with the proposed HetGAT model.
In the training, the same hyperparameters used in the previous experiments are used here, as well. Figure 6 presents the training history for the three models, and shows that GCN and GAT models struggle during the training process. In contrast, the HetGAT model can successfully learn the flow dynamics across multiple graphs, leading to a decrease in training loss over time. For a graph size of 100, GCN and GAT had a training loss of 0.114 and 0.124, respectively, whereas HetGAT achieved a training loss of 0.032, representing a 70% improvement. The prediction performance metrics on testing sets are compared in Table 4, where the mean absolute error of HetGAT outperformed GAT and GCN for graph sizes of 100 and 300, with values of 0.033 and 0.051, respectively.
(a) Graph Size of 100
(b) Graph Size of 300
Figure 6: Training loss history under generalized synthetic networks. Two network sizes of 100 and 300 are selected for training and testing. Two benchmarks including GCN and GAT are compared with HetGAT.
@@
Table 4: Comparison of the performance of HetGAT with that of GAT and GCN on generalized synthetic networks. Two graph sizes of 100 and 300 are considered in the experiments. The mean absolute error (MAE), normalized conservation loss ($L_{f}$), and coefficient of determination ($R^{2}$) are used to evaluate the prediction performance on the testing set.
Graph Size Metric Model
GAT GCN HetGAT
100 MAE 0.143 0.179 0.033
$L_{f}$ 0.016 0.014 0.013
$R^{2}$ 0.602 0.372 0.979
300 MAE 0.291 0.316 0.051
$L_{f}$ 0.018 0.016 0.013
$R^{2}$ 0.412 0.316 0.977
@@ 5 Conclusion and Discussion
In this paper, we proposed a novel approach for traffic assignment and traffic flow learning using heterogeneous graph neural networks. We conducted extensive experiments on three real-world traffic networks to evaluate the performance of the proposed heterogeneous GNN model and compared it with the state-of-the-art models. The results show that the proposed model outperforms other models in terms of convergence rate, training loss, and prediction accuracy. The proposed model can also be generalized to different network topologies, with different topologies or network parameters, which demonstrates its potential in real-world applications. Currently, the proposed HetGAT model only learns and predicts the static traffic flow patterns. As a potential extension of this work, the proposed framework can be extended to learn the dynamic traffic flow patterns. Furthermore, the current proposed GNN model uses training data collected from conventional solvers of static traffic assignment. In future work, we will study how these models can be trained on traffic data collected from sensors, such as loop detectors, cameras, or GPS devices.
6 Acknowledgment

### v2

@@ 4.2 Experiments on Generalized Synthetic Networks
In this section, unlike the experiments in Section 4.1, which involved training and testing on an identical network topology, we examine the generalization capability of the proposed model to networks with varied topologies. In particular, we consider real-world scenarios in which certain links in the network are fully closed due to governmental directives or catastrophic events such as bridge collapses, leading to significant alterations in network topology. Another scenario for topology alternation is when cities consider network expansion to better serve increased mobility demands due to current urbanization trends. Under all these scenarios, the resulting urban networks may exhibit both commonalities and disparities in their topologies (Rodrigue, 2020). This is while training models separately for each distinct network requires substantial time and effort. Motivated by the aforementioned considerations, our aim is to explore the generalization ability of our HetGAT model over varying topologies.
@@
Table 4: Performance comparison of HetGAT with benchmark methods on modified urban transportation network with link addition and removal. The mean absolute error, root mean square error, and normalized conservation loss are used to evaluate the prediction performance on the testing set.
Network Model Missing Ratio = 0% Missing Ratio = 20% Missing Ratio = 40%
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
Sioux GCN 10.12 13.97 8.31% 11.77% 0.15 10.96 14.82 9.22% 13.05% 0.15 10.94 15.04 9.05% 12.86% 0.15
GAT 7.16 9.89 6.10% 8.77% 0.13 10.78 14.40 9.11% 12.77% 0.14 12.20 16.42 10.34% 14.58% 0.17
HetGAT 3.04 4.22 2.56% 3.75% 0.07 3.33 4.53 2.79% 3.96% 0.07 3.46 4.69 2.88% 4.08% 0.08
EMA GCN 4.60 7.82 15.11% 26.42% 0.26 4.09 6.68 13.97% 24.71% 0.24 4.04 6.57 15.24% 31.99% 0.23
GAT 2.70 4.50 9.97% 18.21% 0.14 2.73 4.55 10.04% 20.05% 0.15 2.86 4.72 10.50% 19.94% 0.15
HetGAT 1.28 2.39 4.09% 7.91% 0.06 1.34 2.47 4.29% 8.17% 0.07 1.42 2.57 4.56% 8.42% 0.07
Anaheim GCN 6.02 9.93 16.29% 26.78% 0.40 5.93 9.83 15.98% 26.40% 0.31 6.04 9.99 16.24% 26.73% 0.37
GAT 5.43 9.16 14.94% 25.20% 0.25 5.37 9.02 14.83% 25.01% 0.40 5.53 9.28 15.23% 25.60% 0.34
HetGAT 1.14 2.22 3.29% 6.79% 0.09 1.15 2.08 3.34% 6.49% 0.09 1.29 2.38 3.71% 7.40% 0.10
In this work, we conduct experiments on two sets of networks: (1) modified urban networks of Section 4.1; (2) synthetic networks. The first set, includes networks that are modifications of the original Sioux Falls, EMA, and Anaheim networks by adding and removing links. For each of these networks, we generate 20 unique topological variations. Additionally, we incorporate three configurations of OD demand in our experiments: complete OD setting, 20% and 40% incomplete OD setting. The training setting and the hyperparameters remain the same as those in the aforementioned experiments in Section 4.1. The results are presented in Table 4. It can be seen that compared to the results in Table 2, HetGAT exhibits comparable levels of effectiveness for all networks. Additionally, in comparison to other baseline models, our proposed HetGAT model consistently outperforms both GAT and GCN over all networks, under both complete and incomplete OD scenarios. It is noteworthy that compared to the homogeneous GNN model, since our model has captured the influence of OD pairs via virtual links, the HetGAT model is more expressive and can better capture and learn the impact of topology alterations on flow distributions.
Figure 7: The histogram of the link capacity and normalized OD demand in the training and testing data of synthetic networks. Two sizes of datasets are included: 100 and 300.
The second set includes synthetic networks. We generate these networks by starting with a grid graph and then adding links between randomly selected nodes. Furthermore, to emulate the real road network, a number of nodes and edges are randomly removed until the number of nodes in the graph reaches a predefined threshold. Two sets of synthetic networks are considered: one with 100 nodes and another with 300 nodes. For each graph size, 20 different graph topologies are generated. Three of these randomly generated graphs of size 100 are shown in Figure 8. The OD demand and the link capacity are also randomly generated using the scaling factor according to Equations 15 and 16, respectively. To demonstrate the diversity of these examples, similar to Section 4.1, the histograms of the link capacities and the OD demands in the generated dataset are shown in Figure 7, indicating the training and testing data are sufficiently diverse to cover different scenarios.
In this work, we consider a case in which the training and test graphs are of the same size, and a case where the training and test graphs are of different size (e.g., training graphs of size 100 and test graphs of size 300). It should be noted that the proposed HetGAT involves a node feature whose dimension is equal to the total number of nodes in the graph. Therefore, this requires an adjustment when training and testing are done on graphs with different number of nodes. In additional to standard training, which trains and tests on the same set of graphs, we propose two training strategies, enabling the HetGAT model to be trained and tested effectively on various graphs:
  • 1.
    Transfer learning: we adapt the model trained from previous experiments by substituting the preprocessing and final layers with new parameters, while freezing the remaining parameters unchanged. This facilitates a quick adaption to the new topologies with limited training.
  • 2.
    Homogenized training: we define a maximum node number threshold $N_{\max}$ and for all the graphs graphs containing fewer nodes than this threshold, we introduce dummy nodes to the graph until the number of node reaches the $N_{\max}$. This approach ensures uniform formulation among different graphs, allowing a single model to train and test across graphs of varying sizes.
Figure 8: The illustrations of sampled generalized synthetic networks with the network size of 100. The link color represents the link capacity of each link.
The training setting and the hyperparameters remain the same as those in the previous experiments. The prediction performance metrics for the testing sets are presented in Table 5. Compared with transfer learning and homogenized learning settings, the proposed model achieves the highest accuracy when same graph size is used in training and testing. Despite a performance decline under the transfer and homogenized learning setting, the HetGAT model maintains a competitive edge, outperforming baseline models such as GAT and GCN.
Table 5: Comparison of the performance of HetGAT with that of GAT and GCN on generalized synthetic networks. Two graph sizes of 100 and 300 are considered in the experiments. We consider three different training strategies: standard training, transfer learning, and homogenized learning.
Network Model Standard training Transfer learning Homogenized learning
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
100 GCN 1.62 2.40 17.85% 26.89% 0.38 1.82 2.63 20.09% 29.70% 0.10 1.45 2.08 15.76% 23.11% 0.11
GAT 1.68 2.48 18.40% 27.49% 0.33 1.92 2.79 21.11% 31.30% 0.10 1.63 2.42 17.90% 27.59% 0.11
HetGAT 0.25 0.38 2.75% 4.27% 0.19 0.38 0.56 4.18% 6.29% 0.09 0.74 1.16 8.46% 14.35% 0.09
300 GCN 2.97 4.11 31.87% 44.14% 0.15 3.09 4.28 33.17% 45.97% 0.13 3.04 4.23 32.65% 45.32% 0.14
GAT 2.81 3.89 30.10% 41.58% 0.18 2.94 4.05 31.43% 43.33% 0.15 2.85 3.94 30.56% 42.19% 0.16
HetGAT 0.46 0.69 4.99% 7.51% 0.11 0.78 1.17 8.45% 12.89% 0.06 0.52 0.77 5.63% 8.45% 0.13
@@ 5 Conclusion and Discussion
In this paper, we proposed a novel approach for end-to-end traffic assignment and traffic flow learning using heterogeneous graph neural networks. Compared with conventional homogeneous graph neural networks, we introduce the auxiliary virtual links connecting origin-destination node pairs. Furthermore, we proposed a novel adaptive graph attention mechanism to effectively capture the semantic and contextual features through different types of links. Furthermore, we conducted extensive experiments on three real-world urban transportation networks to evaluate the performance of the proposed model and compared it with the state-of-the-art models. The results indicate that the proposed model outperforms other models in terms of convergence rate and prediction accuracy. Notably, by introducing two different training strategies, the proposed heterogeneous graph neural network model can also be generalized to different network topologies, underscoring its potential in real-world scenarios.
Our work can be considered as an efficient surrogate model for further accelerating the complex optimization tasks including resource allocation and infrastructure asset management. Currently, the proposed HetGAT model only learns and predicts the static traffic flow patterns. As a potential extension of this work, the proposed framework can be extended to learn the dynamic traffic flow patterns. Furthermore, the current proposed GNN model uses training data collected from conventional solvers of static traffic assignment. In future work, we will explore how these models can be trained on traffic data collected from sensors, such as loop detectors, cameras, or GPS devices.
Acknowledgment